Paul Coddington SS No: 604-16-3234 Northeast Parallel Architectures Center 111 College Place, 3-217 CST Syracuse University Email: paulc@npac.syr.edu

April 17, 1996

Payroll Office Skytop Building Syracuse University

To whom it may concern,

Please change my home and work address in your records. Starting immediately, my new address will be

Home: 226 Degraw St, Apt 1

Brooklyn NY 11231 Tel: (718) 243-2413

Work: Lubin House

11 E. 61st St New York NY 10021 Tel: (212) 826-0333 ext 140

Yours sincerely,

Paul Coddington NPAC Research Scientist